OpenStudy (anonymous):

at what points are the tangents to the graph of f(x)=x^2-3x horizontal?

OpenStudy (anonymous):

set the derivative equal to zero and solve

OpenStudy (anonymous):

\[f(x)=2x^2-3x\] \[f'(x)=2x-3\] set it equal to zero and solve you don't really need calc for this. this is a quadratic, and slope is zero at the vertex first coordinate of the vertex is \(-\frac{b}{2a}\) which is what you get when you take the derivative, set equal zero and solve

OpenStudy (anonymous):

ok so i get (-3/2)... im also supposed to somehow get the point (-9/4). how do i do that?

OpenStudy (anonymous):

you have the function right?

OpenStudy (anonymous):

it is \((\frac{3}{2},f(\frac{3}{2}))\)

OpenStudy (anonymous):

oh btw it is \[2x-3=0\implies x=\frac{3}{2}\] not negative

OpenStudy (anonymous):

oh sorry thats what i meant! and i get it! thanksss!

OpenStudy (anonymous):

yw